How old is Charlotte if her age now is equal to 3 times her age in 3 years, less 3 times her age 3 years ago?

Answer by htmentor(1343)

Let x = Charlotte's current age
Her age now (x) is equal to 3 times her age in 3 years (x+3) less 3 times her age 3 years ago (x-3). In equation form this is:
x = 3(x+3) - 3(x-3)
Solve for x:
x = 3x + 9 - 3x + 9
x = 18
Check:
3*21 - 3*15 = 63 - 45 = 18
Ans: 18